# Tethering Broken Themes: Aligning Neural Topic Models with Labels and Authors

### v1

@@ Appendix A Modeling Details
All of the aforementioned models are also constructed with their corresponding FANToM variants. In these variants, only the priors are manipulated in the objective function (as mentioned in described in Section 4.2 of the main paper), and a separate author decoder is introduced to learn the topic-author multinomial distribution. The core architecture remains unchanged in all cases.
@@
word embeddings (Pennington et al., 2014, GloVe)
train:val:test 70:15:15
Appendix B FANToM
Figure 7: The graphical model of our framework is also shown, where solid lines represent the generative distribution $p$, and dotted lines represent the variational distribution $q$.
B.1 Illustrative example to construct topic-label vector
[…]
Using this label indicator vector, the modified prior parameter is defined as $\gamma=\alpha\cdot\mathbb{I}^{\text{d}}$, where $\alpha$ is a scaling factor. This allows for flexible topic assignments based on the available labels. For example, if $L_{k}=$ (sport, sport, cars, cars, weather, weather, no-label, no-label) , multiple topics may be linked to a single label. This allows the model to learn multiple focused topics.
B.2 Training FANToM
The training algorithm for FANToM, outlined in Algorithm 1, begins by taking as input a document set $D$, an expert $E$, a prior parameter $\alpha$, and topic-label assignments $L_{k}$. The process starts with the expert providing labels $l$ for the documents, which are then used to create a multi-hot vector $\mathbb{I}^{\text{d}}$ that reflects the association between labels and topics.
The algorithm processes batches $\mathcal{B}$ of documents from the set $D$. For each batch, the bag-of-words (BoW) representation $x$ and an author representation $a$ are extracted. An encoder network, parameterized by $\theta$, computes the posterior parameters $\alpha_{p}$ from the BoW representation. Simultaneously, a prior distribution is constructed using the expert-provided labels, with parameters given by $\gamma=\alpha\cdot\mathbb{I}^{\text{d}}$.
A variable $z$ is then sampled from a Dirichlet distribution $\text{Dirichlet}(\alpha_{p})$, representing the document-topic distribution. This sampled $z$ is input into two decoder networks: one reconstructs the document as $x^{\prime}$, while the other reconstructs the author representation as $a^{\prime}$.
The model parameters are updated by minimizing the objective function defined in Eq. 2. This iterative process of batch processing, parameter updating, and reconstruction continues until the algorithm converges, ensuring the model effectively learns both document and author representations aligned with the provided topic labels.
Throughout the training, the interaction between the posterior parameters derived from the encoder, the prior informed by expert labels, and the outputs from the decoders continually refines the model’s understanding of topic, label, and author associations. This optimization process ensures that FANToM captures the underlying topic structures in the document set $D$, aligning them with expert-provided labels and authors.
Algorithm 1 Training FANToM
1:  Input: Documents $D$, Expert $E$, Prior parameter $\alpha$, Topic-label assignments $L_{k}$
2:  $\lambda\leftarrow$ Get labels from expert $E$ for documents $D$
3:  $\mathbb{I}^{\text{d}}\leftarrow$ Create multi-hot vector based on $\lambda^{\text{d}}$ and $L_{k}$ for $d\in D$
4:  Initialize model parameters $(\theta,\phi,\psi)$
5:  while not converged do
6:     for batch $\mathcal{B}$ in $D$ do
7:        Extract BoW $x$, authors $a$, and multi-hot vector $\mathbb{I}^{\text{d}}$ from $\mathcal{B}$
8:        $\gamma\leftarrow\alpha\cdot\mathbb{I}^{\text{d}}$ (modify prior parameter)
9:        $\alpha_{p}\leftarrow\text{Encoder}(\theta;x)$ (encode words)
10:        Sample $z\sim\text{Dirichlet}(\alpha_{p})$
11:        $x^{\prime}\leftarrow\text{Decoder}(\phi;z)$ (reconstruct document)
12:        $a^{\prime}\leftarrow\text{Decoder}(\psi;z)$ (reconstruct author)
13:        Compute gradients $\nabla_{(\theta,\phi,\psi)}L_{\text{F}}(\theta,\phi,\psi;x,\gamma)$ (Eq. 2)
14:        Update parameters $(\theta,\phi,\psi)$
15:     end for
16:  end while
Appendix C Datasets and Preprocessing
@@
For dataset preparation, we use SpaCy Honnibal and Montani (2017) for data tokenization. Additionally, we eliminate common stop words, punctuations, as well as high and low-frequency words. High and low-frequency words are determined by excluding words that appear in over 85% of the documents or in fewer than 30 documents. This standardization is applied across all models.
Appendix D Using LLMs as Expert
@@
Not all datasets come with accompanying labels and labeling every corpus can be a challenging task. However, an alternative approach involves using Large Language Models (LLMs) Ostheimer et al. (2023, 2024) to assign labels to unlabeled documents. In our study, we explore this possibility by using the capabilities of BART Lewis et al. (2019) for zero-shot text classification. To achieve this, we use a pretrained model from Wolf et al. (2019), which undergoes fine-tuning on Natural Language Inference (NLI) Williams et al. (2018) following the methodology described by Yin et al. (2019).
@@
This method operates by treating the sequence to be classified as the NLI premise and constructing a hypothesis for each potential label. For example, when determining if a sequence pertains to the "sport" label, we can formulate a hypothesis such as "This text is about sports." By converting the probabilities for entailment and contradiction, we obtain the label probabilities.
@@ Appendix E Runtime
Appendix F Additional Evaluation
@@ F.1 Additional Quantitative Evaluation
Tables 9 and 10 delve into topic coherence and diversity in comparison to baselines, thus demonstrating the superior quality of topics learned by our model.
We also validate alignment quantitatively by conducting text classification using FANToM(L). We use the encoder to extract posterior probabilities for the test set during inference. We evaluate each model based on Top-3 and Top-5 accuracy and F1 scores. As demonstrated in Tables 7 and 8, FANToM(L) yields good classification and F1 scores which suggests alignment with the expert-assigned labels.
We also conduct benchmarking experiments in a semi-supervised scenario where not all topics in the corpus have labels. To manage this, we designate the "no-label" category for documents lacking labels. In our experiment, we randomly remove labels from 50% of the documents across all datasets and train models to learn 50 and 200 topics. This setup results in a mixture of labeled and "no-label" topics, reflecting real-world situations where not all documents may have known labels. Analyzing the results, shown in Table 4, FANToM consistently outperforms the baseline models, demonstrating its robustness even in the absence of complete label information. Specifically, the FANToM variants show higher topic coherence (TC) and topic quality (TQ) across datasets. Despite the removal of labels, FANToM maintains superior peformance and consistently achieves better overall scores, solidifying its advantage in semi-supervised settings.
F.2 Topic-Author Alignment
Our framework allows us to link topics with authors using two distinct approaches: FANToM, which uses labels, and FANToM(A), which does not. In Table 5, we present results that demonstrates that our method associates topics with authors and labels from the arxiv dataset. The authors associated with the topics belong to their respective field of research. Including labels allows us to validate this association by checking the respective topic authors within their labels. Further we also list results from FANToM(A) on stats.ML subset from the arxiv dataset in Figure 11.
Appendix G Detailed Discussion on Topic Models and LLMs
In recent years, combining topic modeling with large language models (LLMs) has emerged as a key area of research, aiming to harness the strengths of both techniques.
Traditionally, topic modeling has employed methods like Latent Dirichlet Allocation (LDA) and its variants to uncover hidden topics in a corpus based on word co-occurrence patterns. However, these methods often face challenges with interpretability and coherence.
The introduction of Variational Autoencoders (VAEs) brought a new level of flexibility to topic modeling by incorporating advancements in natural language processing (NLP), such as word embeddings. Later, with the advent of transformer models like BERT, contextualized embeddings were integrated into topic modeling approaches, such as in Contextualized Topic Models (CTM).
The rise of LLMs, known for their ability to generate coherent and contextually relevant text, has opened up new possibilities for enhancing topic modeling. One notable development is the use of prompt engineering, exemplified by recent approaches like TopicGPT Pham et al. (2024). This method involves prompting an LLM to generate topics by processing each document individually. While this approach can produce high-quality topics, it incurs substantial computational costs due to the need to process prompts for every document, which can be particularly demanding with large corpora and complex LLMs. Additionally, prompting methods often lack a systematic framework for incorporating information such as authors and labels into the topic modeling process.
FANToM offers an alternative approach by integrating LLMs and topic modeling in a different way. It continues to use a VAE framework, known for its efficiency in topic modeling, while employing a smaller, task-specific LLM for labeling tasks. The VAE handles the core topic modeling process, effectively capturing the latent structure of the data and providing compact and interpretable topic representations. For labeling, FANToM utilizes a smaller, specialized LLM rather than a large-scale general-purpose model, which reduces computational demands while maintaining high accuracy. By aligning the outputs of the smaller LLM with the topics generated by the VAE, FANToM achieves a more coherent and adaptable solution that can be tailored to various datasets and applications.
Overall, FANToM opens up new possibilities for leveraging large-scale LLMs to enhance traditional topic modeling frameworks, leading to improved topic generation and alignment based on specific use cases.
Models 20NG AGN DB-14 arxiv
TC TD TQ TC TD TQ TC TD TQ TC TD TQ
50 Topics
baseline SCHOLAR $0.410$ $0.57$ $0.234$ $0.372$ $0.44$ $0.164$ $0.629$ $0.68$ $0.428$ $0.364$ $0.46$ $0.168$
DVAE $0.354$ $0.59$ $0.209$ $0.393$ $0.46$ $0.181$ $0.658$ $0.66$ $0.434$ $0.360$ $0.48$ $0.173$
ETM $0.418$ $0.56$ $0.234$ $0.326$ $0.43$ $0.140$ $0.630$ $0.67$ $0.422$ $0.281$ $0.45$ $0.126$
CTM $0.412$ $0.58$ $0.239$ $0.371$ $0.44$ $0.163$ $0.632$ $0.69$ $0.436$ $0.348$ $0.47$ $0.164$
ECRTM $0.465$ $0.60$ $0.279$ $0.389$ $0.45$ $0.175$ $0.670$ $0.71$ $0.475$ $0.390$ $0.49$ $0.191$
FANToM (ours) SCHOLAR $0.415$ $0.62$ $0.257$ $0.397$ $0.48$ $0.191$ $0.608$ $0.72$ $0.438$ $0.375$ $0.50$ $0.188$
DVAE $0.424$ $0.61$ $0.258$ $0.409$ $0.50$ $0.205$ $0.611$ $0.71$ $0.434$ $0.336$ $0.51$ $0.171$
ETM $0.410$ $0.60$ $0.246$ $0.391$ $0.49$ $0.191$ $0.639$ $0.70$ $0.447$ $0.310$ $0.49$ $0.151$
CTM $0.436$ $0.64$ $0.279$ $0.395$ $0.47$ $0.186$ $0.622$ $0.69$ $0.429$ $0.342$ $0.52$ $0.178$
ECRTM $0.440$ $0.65$ $0.286$ $0.400$ $0.50$ $0.200$ $0.645$ $0.73$ $0.471$ $0.395$ $0.53$ $0.209$
200 Topics
baseline SCHOLAR $0.362$ $0.47$ $0.170$ $0.354$ $0.38$ $0.134$ $0.581$ $0.57$ $0.331$ $0.342$ $0.41$ $0.140$
DVAE $0.376$ $0.49$ $0.184$ $0.374$ $0.39$ $0.146$ $0.602$ $0.58$ $0.349$ $0.335$ $0.43$ $0.144$
ETM $0.368$ $0.48$ $0.177$ $0.361$ $0.37$ $0.134$ $0.590$ $0.59$ $0.348$ $0.328$ $0.40$ $0.131$
CTM $0.384$ $0.50$ $0.192$ $0.368$ $0.40$ $0.147$ $0.595$ $0.60$ $0.357$ $0.340$ $0.42$ $0.143$
ECRTM $0.397$ $0.52$ $0.206$ $0.370$ $0.41$ $0.151$ $0.612$ $0.61$ $0.373$ $0.355$ $0.44$ $0.156$
FANToM (ours) SCHOLAR $0.380$ $0.54$ $0.205$ $0.362$ $0.40$ $0.145$ $0.615$ $0.62$ $0.381$ $0.355$ $0.46$ $0.163$
DVAE $0.387$ $0.53$ $0.205$ $0.381$ $0.42$ $0.160$ $0.617$ $0.61$ $0.376$ $0.360$ $0.48$ $0.173$
ETM $0.379$ $0.52$ $0.197$ $0.385$ $0.41$ $0.158$ $0.620$ $0.60$ $0.372$ $0.365$ $0.45$ $0.164$
CTM $0.399$ $0.55$ $0.220$ $0.380$ $0.39$ $0.148$ $0.624$ $0.63$ $0.393$ $0.369$ $0.49$ $0.181$
ECRTM $0.404$ $0.56$ $0.226$ $0.386$ $0.43$ $0.166$ $0.630$ $0.64$ $0.403$ $0.371$ $0.50$ $0.185$
Table 4: Topic Coherence (TC), Topic Diversity (TD), and Topic Quality (TQ) for baseline and FANToM models across datasets for 50 and 200 topics in semi-supervised setting. In general, FANToM outperforms baseline models across both 50 and 200 topics.
Index/Label Aspect Topics/Authors
2 topic graphs, chromatic, graph, bipartite, isomorphic
authors Kráľ_D, Seymour_P, Alon_N, Wood_DR, Koolen_JH
4 topic exoplanet, orbiting, planet, kepler, comet
authors Ford Eric B., Agol E., Wright J.T., Pepe F., Henning T.
math.CO topic conjectured, automorphism, matroids, poset, extremal
authors Seymour_P, Kráľ_D, Rautenbach_D, Klavžar_S, Li_X
astro-ph.EP topic stellar, atmospheric, exoplanet, atmosphere, planet, orbiting
authors Henning T., Lagrange A. -M, Pepe F., Desidera S., Ford Eric B.
Table 5: Associations between Topics and Authors using FANToM(A) (top) and FANToM (bottom) approaches. FANToM builds meaningful correspondence between labels, topics, and authors.
@@ 50 Topics
Figure 10: This figure compares FANToM(L) (top), Labeled LDA (middle), and Neural Topic Models (bottom) on the 20NG dataset. FANToM(L) exhibits aligned and esoteric topics consistent with human intentions, while Labeled LDA shows repetitive generic words in many topics (e.g., "host," "writes," "articles"), leading to low diversity. Neural Topic Models do not have association with provided labels and fail to produce certain topics completely, unlike FANToM(L) (e.g., "mac", "graphics", "electronics").
Figure 11: Topics and their top authors from "stats.ML" label in the arxiv dataset between 2010-2020. The authors are correctly associated with their research area.

### v2

# Appendix for FANToM: A Framework to Align Neural Topic Models

@@ Appendix A Modeling Details
All of the aforementioned models are also constructed with their corresponding FANToM variants. In these variants, only the priors are manipulated in the objective function (as mentioned in described in Section LABEL:align-topic-label of the main paper), and a separate author decoder is introduced to learn the topic-author multinomial distribution. The core architecture remains unchanged in all cases.
@@
word embeddings [pennington2014glove, GloVe]
Appendix B Datasets and Preprocessing
@@
For dataset preparation, we use SpaCy [spacy2] for data tokenization. Additionally, we eliminate common stop words, punctuations, as well as high and low-frequency words. High and low-frequency words are determined by excluding words that appear in over 85% of the documents or in fewer than 30 documents. This standardization is applied across all models.
Appendix C Using LLMs as Expert
@@
Not all datasets come with accompanying labels and labeling every corpus can be a challenging task. However, an alternative approach involves using Large Language Models (LLMs) to assign labels to unlabeled documents. In our study, we explore this possibility by using the capabilities of BART [lewis2019bart] for zero-shot text classification. To achieve this, we use a pretrained model from [wolf2019huggingface], which undergoes fine-tuning on Natural Language Inference (NLI) [N18-1101] following the methodology described by [yin2019benchmarking].
@@
This method operates by treating the sequence to be classified as the NLI premise and constructing a hypothesis for each potential label. For example, when determining if a sequence pertains to the ”sport” category, we can formulate a hypothesis such as ”This text is about sports.” By converting the probabilities for entailment and contradiction, we obtain the label probabilities.
@@ Appendix D Runtime
Appendix E Additional Evaluation
@@
In addition to qualitative findings, we introduce supplementary quantitative results. Table 4 presents Top-3 and Top-5 accuracy scores of the FANToM(L) model validates the alignment of topics with labels. Table 3 presents F1 scores, affirming the successful alignment. Tables 5 and 6 delve into topic coherence and diversity in comparison to baselines, thus demonstrating the superior quality of topics learned by our model.
@@
Figure 3: This figure compares FANToM(L) (top), Labeled LDA (middle), and Neural Topic Models (bottom) on the 20NG dataset. FANToM(L) exhibits aligned and esoteric topics consistent with human intentions, while Labeled LDA shows repetitive generic words in many topics (e.g., ”host,” ”writes,” ”articles”), leading to low diversity. Neural Topic Models do not have association with provided labels and fail to produce certain topics completely, unlike FANToM(L) (e.g., ”mac”, ”graphics”, ”electronics”).
Figure 4: Topics and their top authors from ”stats.ML” category in the arxiv dataset between 2010-2020. The authors are correctly associated with their research area.
@@
Appendix F Limitations
One common limitation in the field of topic models is the reliance on automated evaluation measures, including the topic quality measure. While topic coherence, a component of the topic quality measure, has demonstrated a reasonable correlation with human evaluations, it is important to note that qualitative assessment of topics remains subjective, even in human evaluations. Consequently, even with human studies, there is no guarantee of consistent and reproducible results. However, it is worth noting that this challenge is pervasive across the field and not specific to our proposed models. Despite these limitations, our proposed models have consistently performed on par with or outperformed existing state-of-the-art models.
As a second limitation, we rely on labels to align our models. The label quality may vary between datasets, influencing the final results. We have not investigated the quality of the labelings, the human annotations in the datasets, or the LLM-assigned labels. Therefore, the results may be slightly biased depending on the quality of the provided labels.
Appendix G Ethical Statement
In this study, we incorporated authorship information into topic models while ensuring ethical considerations. To acquire the dataset, we obtained it from the openly available [arxiv], which is licensed under the CC0 1.0 Universal (CC0 1.0) Public Domain Dedication, allowing us unrestricted use of the data. We specifically utilized the abstracts and author names of openly available papers. It is important to note that we have mentioned author names as they were presented to us by our model, without making any alterations. We did not have access to any sensitive or restricted information.